# Can LLMs Simulate Human Behavioral Variability? A Case Study in the Phonemic Fluency Task

### v1

@@ 6 Discussion
This study focuses on a single verbal fluency task (letter F) in English, limiting generalizability across task types, letters, and languages. While phonemic fluency provides a stringent test of form-based retrieval, different letters may elicit distinct lexical or structural patterns. Similarly, results may differ in languages with different orthographic or phonological systems. Our analysis also centers on a specific subset of models and parameter settings; other prompting techniques (e.g., few-shot examples, chain-of-thought reasoning) or fine-tuned models may yield different patterns. Additionally, while we simulated participant-level responses using demographic and performance metadata, we did not incorporate richer individual-level traits (e.g., reading history, cognitive profiles) that may be necessary to drive human-like variability. Future work should expand to multilingual and cross-task comparisons, explore more cognitively grounded persona modeling, and evaluate whether newer architectures or sampling strategies better capture human behavioral diversity.
Appendix A Overview of LLM Providers, Prompts, and Sampling Temperatures
We evaluated 34 model configurations across six providers–OpenAI, Anthropic, Google, xAI, Meta (LLaMA), and Alibaba (Qwen)–systematically varying model family (including general-purpose vs. reasoning-focused models and different model generations), temperature, and prompt specificity. In subsequent analyses, we also examined ensemble combinations of these models to assess response variability. Table A1 summarizes these configurations. Prompt specificity ranged from full participant profiles to minimal or absent information, designed to test how different inputs affected response count and variability.
Provider | Model | Temperature | Prompt Type |
OpenAI (15 model configurations) | |||
gpt-4.1-2025-04-14 | 0.3, 0.7 (default), 1.1, 1.5 | Full (all temperatures); No Info, Demographic, Performance (0.7 only) | |
gpt-4o-2024-08-06 | 0.7 (default) | Full | |
gpt-4-turbo-2024-04-09 | 0.7 (default) | Full | |
gpt-3.5-turbo-0125 | 0.7 (default) | Full | |
[…]
o3-2025-04-16 (reasoning) | – | Full | |
o3-mini-2025-01-31 (reasoning) | – | Full | |
Anthropic (8 model configurations) | |||
claude-3-7-sonnet-20250219 | 0.3, 0.7, 1.0 (default) | Full (all temperatures); No Info, Demographic, Performance (0.7 only) | |
claude-3-5-sonnet-20241022 | 1.0 (default) | Full | |
[…]
Google (3 model configurations) | |||
gemini-2.5-flash-preview-04-17 | default | Full | |
gemini-2.0-flash | default | Full | |
[…]
xAI (4 model configurations) | |||
grok-3-latest | 0.7, 1.0 | Full | |
[…]
Open-Source Models (using together.ai API; 4 model configurations) | |||
meta-llama/Llama-4-Maverick-17B-128E-Instruct-FP8 | default | Full | |
meta-llama/Llama-4-Scout-17B-16E-Instruct | default | Full | |
meta-llama/Llama-3.3-70B-Instruct-Turbo | default | Full | |
[…]
Prompt types: Full = complete demographic and performance information; Demographic = only demographic information (age, education); No-Info = no participant information; Performance = only performance information.
@@ Appendix A Overview of LLM Providers, Prompts, and Sampling Temperatures
Appendix B Stepwise Regression Analyses of Linguistic Predictors of Production Frequency
@@
Among the ensemble models, the Top 5 Models ensemble yielded a significant model [] explaining of variance (adjusted ), with age of acquisition as the primary predictor () followed by word frequency (). The 100+ Word Types ensemble produced a one-predictor model [] with of variance explained (adjusted ), with age of acquisition as the sole predictor (). The All-Model Mix similarly retained only age of acquisition [], explaining of variance (adjusted , ).

### v2

@@ 7. Discussion and Conclusion
This study focuses on a single verbal fluency task (letter F) in English, limiting generalizability across task types, letters, and languages. While phonemic fluency provides a stringent test of form-based retrieval, different letters may elicit distinct lexical or structural patterns. Similarly, results may differ in languages with different orthographic or phonological systems. Our analysis also centers on a specific set of models and parameter settings; other prompting techniques (e.g., few-shot examples) or fine-tuned models may yield different patterns. Additionally, while we grounded our simulations in real participant metadata, age and education are not the only sources of individual variation in verbal fluency. Factors such as profession, domain expertise, and accumulated life experience also contribute to behavioral differences, and incorporating these richer individual-level traits may be necessary to drive more human-like variability. Future work should expand to multilingual and cross-task comparisons and evaluate whether alternative sampling or persona modeling strategies better capture human behavioral diversity.
Ethical Considerations
The human data used in this study are drawn from Qiu and Johns (2021), a publicly available dataset under a CC-BY 4.0 license. The dataset includes anonymized verbal fluency responses alongside participant age and education information, with all personal identifiers removed. According to the original publication, all data collection procedures received institutional ethics approval, and participants provided informed consent. In our study, only de-identified demographic and performance information was used to construct LLM prompts. At no stage do we include real human response content in the prompts, nor do we attempt to reconstruct the linguistic patterns of specific individuals.
Our findings also raise ethical concerns regarding the use of LLMs to simulate human behavioral variability. LLMs exhibit output rigidity, favoring high-frequency, prototypical responses and failing to capture natural human diversity. If used as substitutes for human participants without acknowledging this limitation, LLM-based simulations risk systematically underrepresenting the variability present in real populations, particularly marginalizing rare or idiosyncratic responses. We therefore caution against treating LLM-simulated individuals as direct substitutes for real human participants, especially in contexts where diversity or individual identity are central to the research question.
Bibliographical References
- The English lexicon project. Behavior Research Methods 39 (3), pp. 445–459. External Links: Document Cited by: §4.2.
- When pumpkin is closer to onion than to squash: The structure of the second language lexicon. Cognition 156, pp. 60–70. External Links: Document Cited by: §5.1, §5.1, §5.2.
- Category norms with a cross-sectional sample of adults in the united states: consideration of cohort, age, and historical effects on semantic categories. Behavior Research Methods 53, pp. 898–917. External Links: Document Cited by: §3.2.
- Automated analysis of digitized letter fluency data. Frontiers in Psychology 12, pp. 654214. External Links: Document Cited by: §1, §4.2.
- The limitations of large language models for understanding human language and cognition. Open Mind 8, pp. 1058–1083. External Links: Document Cited by: §1.
- Can ai language models replace human participants?. Trends in Cognitive Sciences 27 (7), pp. 597–600. External Links: Document Cited by: §1.
- On the evolution of random graphs. Publ. Math. Inst. Hung. Acad. Sci 5 (1), pp. 17–60. Cited by: §5.2.
- Norms for letter and category fluency: demographic corrections for age, education, and ethnicity. Assessment 6 (2), pp. 147–178. External Links: Document Cited by: Appendix A, §1.
- Bias in large language models: origin, evaluation, and mitigation. arXiv preprint arXiv:2411.10915. External Links: Document Cited by: §7.
- Cognitive networks highlight differences and similarities in the stem mindsets of human and llm-simulated trainees, experts and academics. arXiv. External Links: Document Cited by: §1.
- Semantic features of object concepts generated with gpt-3. In Proceedings of the Annual Meeting of the Cognitive Science Society, Vol. 44. External Links: Link Cited by: §1.
- More agents is all you need. arXiv. External Links: Document Cited by: §6.
- A network analysis of the semantic evolution of ‘fruit’ and ‘stone’ in Tibeto-Burman languages. Poznan Studies in Contemporary Linguistics 61 (2), pp. 121–172. External Links: Document Cited by: §5.1.
- Network filtering for big data: Triangulated maximally filtered graph. Journal of Complex Networks 5 (2), pp. 161–178. External Links: Document Cited by: §5.1.
- A toolbox for representational similarity analysis. PLoS Computational Biology 10 (4), pp. e1003553. External Links: Document Cited by: §5.2.
- Structural comparisons of noun and verb networks in the mental lexicon. In Proceedings of the Annual Meeting of the Cognitive Science Society, Vol. 43, pp. 1649–1655. External Links: Link Cited by: §2.1, §5.2.
- Estimating type of print exposure across aging through author production. In Proceedings of the Annual Meeting of the Cognitive Science Society, Vol. 46, pp. 1175–1181. External Links: Link Cited by: §7.
- A distributional and sensorimotor analysis of noun and verb fluency. PsyArXiv. External Links: Document Cited by: §2.1, Ethical Considerations.
- Investigating the network structure of domain-specific knowledge using the semantic fluency task. Memory & Cognition 51 (3), pp. 623–646. External Links: Document Cited by: §5.1.
- Cognitive network science: A review of research on cognition through the lens of network representations, processes, and dynamics. Complexity 2019. External Links: Document Cited by: §5.2.
- A large-scale semantic analysis of verbal fluency across the aging spectrum: Data from the Canadian Longitudinal Study on Aging. The Journals of Gerontology Series B: Psychological Sciences and Social Sciences 75 (9), pp. e221–e230. External Links: Document Cited by: §4.1, §4.2.
- Normative data stratified by age and education for two measures of verbal fluency: FAS and animal naming. Archives of Clinical Neuropsychology 14 (2), pp. 167–177. External Links: Document Cited by: Appendix A, §1.
- Large language models and the wisdom of small crowds. Open Mind 8, pp. 723–738. External Links: Document Cited by: §1.
- A systematic review of normative data for verbal fluency test in different languages. Neuropsychology Review 33 (4), pp. 733–764. External Links: Document Cited by: Appendix A, §1.
- Can llm simulations truly reflect humanity? a deep dive. In ICLR Blogposts 2025, External Links: Link Cited by: §1, §7.
- The fluency-based semantic network of LLMs differs from humans. Computers in Human Behavior: Artificial Humans 3, pp. 100103. External Links: Document Cited by: Appendix A, §1, §1, §1.
- Human variability vs. machine consistency: a linguistic analysis of texts generated by humans and large language models. arXiv. External Links: Document Cited by: §1.
Appendix A Prompt Design
Figure A1 shows the complete prompt used to simulate the F fluency task. The prompt included three pieces of participant-level information: age, education level, and number of correct responses.
Age and education were included because they are among the most reliable predictors of verbal fluency performance, consistently accounting for significant variance in both phonemic and semantic fluency across languages and populations (Tombaugh et al., 1999; Gladsjo et al., 1999; Villalobos et al., 2023). These variables were drawn directly from the human dataset rather than fabricated, ensuring ecological validity in the simulation.
In terms of the inclusion of the number of correct responses, Wang et al. (2025b) found that LLMs produced comparable numbers of animal names to human participants when prompted with only a one-minute time constraint and no performance information, suggesting that LLMs can approximate typical human production rates for semantically driven tasks. However, in our pilot testing of the phonemic fluency task, we found that prompting models with only age and education information led to significant overproduction. This likely reflects the greater difficulty of phonemic fluency for humans, who must perform an effortful, non-semantic search through the mental lexicon—a constraint that does not apply to LLMs. To ensure realistic response counts and enable meaningful comparisons at the item and network levels, we therefore also provided the number of correct responses as an additional constraint in the prompt.
Appendix B Model Selection Criteria
Tables A1 and A2 summarize the performance of all 45 configurations in terms of adherence to the instructed number of responses. Mean Absolute Error (MAE) was used to evaluate each configuration across all simulations, capturing the average magnitude of deviation between LLM-generated and human response counts regardless of direction:
where is the number of participants and is the number of correct responses for participant . Configurations with (i.e., 10% of the human mean of ) were considered to have successfully adhered to the performance constraints.
We additionally report Mean Bias Error (MBE), which captures the direction of deviation:
Positive MBE values indicate systematic overproduction relative to human participants, while negative values indicate underproduction. Because LLM-generated responses were subjected to the same error-checking criteria applied to human data—with repetitions being the most common error type when errors occurred, though overall error rates were very low—many models that closely followed the instructed number of responses tended to show slightly negative MBE values after error removal.
Table A1 also reports the number of outlier simulations per configuration, defined as cases where . Of the configurations meeting the MAE criterion, produced no outliers and were retained for participant-level and item-level analyses. The remaining configurations with outliers were excluded from these analyses due to unreliable individual simulations. However, for the LLM ensemble analysis, these configurations were included after removing their outlier simulations, in order to maximize the diversity of LLMs contributing to the ensemble.
| Provider | Model | MAE | MBE | Outliers |
| OpenAI | GPT-5 | 0.00 | 0.00 | 0 |
| GPT-5 mini | 0.00 | 0.00 | 0 | |
| GPT-5.2 (High Reasoning) | 0.00 | 0.00 | 0 | |
| GPT-5.2 (Medium Reasoning) | 0.00 | 0.00 | 0 | |
| GPT-5.2 (Low Reasoning) | 0.00 | 0.00 | 0 | |
| o3-mini | 0.00 | 0.00 | 0 | |
| GPT-4.1 | 0.06 | 0.06 | 0 | |
| GPT-3.5 Turbo | 0.07 | 0.07 | 0 | |
| GPT-4o | 0.16 | 0.16 | 0 | |
| GPT-5.2 | 1.50 | 1.46 | 2 | |
| Anthropic | Claude 4.5 Sonnet (Thinking) | 0.00 | 0.00 | 0 |
| Claude 4 Sonnet (Thinking) | 0.01 | 0.01 | 0 | |
| Claude 3.7 Sonnet (Thinking) | 0.02 | 0.02 | 0 | |
| Claude 4.5 Sonnet | 0.06 | 0.06 | 0 | |
| Claude 3.7 Sonnet | 0.12 | 0.08 | 0 | |
| Claude 4 Sonnet | 0.14 | 0.05 | 0 | |
| Claude 4.6 Sonnet | 0.75 | 0.71 | 3 | |
| Claude 4 Opus | 0.84 | 0.07 | 4 | |
| Claude 4.1 Opus | 1.06 | 0.62 | 6 | |
| Claude 4.5 Opus | 1.43 | 1.30 | 10 | |
| Claude 4.6 Opus (Thinking) | 1.48 | 1.46 | 9 | |
| Gemini 2.5 Pro | 0.00 | 0.00 | 0 | |
| Gemini 3 Pro (High Thinking) | 0.00 | 0.00 | 0 | |
| Gemini 2.5 Flash | 0.04 | 0.04 | 0 | |
| Gemini 3 Pro (Low Thinking) | 0.85 | 0.72 | 2 | |
| Gemini 3 Flash | 0.86 | 0.86 | 8 | |
| xAI | Grok-4.1-fast (Reasoning) | 0.05 | 0.03 | 0 |
| Grok-4-fast (Reasoning) | 0.06 | 0.06 | 1 | |
| Open-Source | Kimi K2 | 0.03 | 0.03 | 0 |
| Qwen3-2507 | 0.08 | 0.08 | 0 | |
| GLM-4.6 | 0.49 | 0.47 | 3 | |
| DeepSeek-V3.1 | 0.97 | 0.93 | 2 | |
| Llama 4 Maverick | 1.29 | 1.27 | 9 |
MAE = Mean Absolute Error; MBE = Mean Bias Error (positive = overproduction, negative = underproduction); Outliers = number of participants with LLMHuman 5.
| Model | MAE | MBE |
| OpenAI | ||
| o4-mini | 3.16 | 3.16 |
| GPT-5.1 | 3.91 | 3.85 |
| o3 | 35.39 | 34.67 |
| GPT-4 Turbo | 57.17 | 57.17 |
| Anthropic | ||
| Claude 4.5 Haiku | 3.12 | 2.92 |
| Claude 4.6 Sonnet (Thinking) | 3.19 | 3.19 |
| Claude 4.6 Opus | 3.87 | 3.81 |
| xAI | ||
| Grok-4-fast | 2.64 | 2.43 |
| Grok-4.1-fast | 3.67 | 3.31 |
| Open-Source Models | ||
| Cogito v2.1 | 3.20 | 2.16 |
| Mistral-Small-24B | 4.05 | 1.48 |
| Ministral-3-14B | 18.62 | 18.62 |
MAE = Mean Absolute Error; MBE = Mean Bias Error.
@@ Appendix B Model Selection Criteria
Appendix C Stepwise Regression Analyses of Linguistic Predictors of Production Frequency
@@
Appendix D Vocabulary Overlap Matrix
Figure A2 presents the pairwise vocabulary overlap coefficients between all LLM configurations retained for participant-level analyses. The overlap coefficient between two vocabularies and is defined as:
A value of indicates that the smaller vocabulary is entirely contained within the larger one, while lower values indicate greater lexical divergence.